# Semi-Supervised Coupled Thin-Plate Spline Model for Rotation Correction and Beyond

### v1

@@ Abstract
Thin-plate spline (TPS) is a principal warp that allows for representing elastic, nonlinear transformation with control point motions. With the increase of control points, the warp becomes increasingly flexible but usually encounters a bottleneck caused by undesired issues, e.g., content distortion. In this paper, we explore generic applications of TPS in single-image-based warping tasks, such as rotation correction, rectangling, and portrait correction. To break this bottleneck, we propose the coupled thin-plate spline model (CoupledTPS), which iteratively couples multiple TPS with limited control points into a more flexible and powerful transformation. Concretely, we first design an iterative search to predict new control points according to the current latent condition. Then, we present the warping flow as a bridge for the coupling of different TPS transformations, effectively eliminating interpolation errors caused by multiple warps. Besides, in light of the laborious annotation cost, we develop a semi-supervised learning scheme to improve warping quality by exploiting unlabeled data. It is formulated through dual transformation between the searched control points of unlabeled data and its graphic augmentation, yielding an implicit correction consistency constraint. Finally, we collect massive unlabeled data to exhibit the benefit of our semi-supervised scheme in rotation correction. Extensive experiments demonstrate the superiority and universality of CoupledTPS over the existing state-of-the-art (SoTA) solutions for rotation correction and beyond. The code and data will be available at https://github.com/nie-lang/CoupledTPS.
@@ 2.1 Rotation Correction
Rigid rotation operation inclines the whole image with a certain angle, but it destroys the rectangular boundaries. To preserve the regular boundaries, He et al. proposed content-aware rotation [14] to rotate the particular contents instead of the whole image. It detected line segments by LSD [15], forced them to rotate to a certain angle, and preserved the contents and boundaries. However, it can only preserve linear structures such as pillars, buildings, etc. When there is a demand to correct a tilted image to the horizontal condition, the specific rotation angle is necessary but unavailable. To acquire this angle prior, whether it is calibrated manually or by automated algorithms [16, 17, 18], the whole process becomes cumbersome and prone to accumulate errors.
To simplify this tedious workflow and free from the constriction of angle prior, Nie et al. proposed a new and practical computer vision task, named rotation correction [10]. It established a rotation correction dataset (DRC-D) with massive triplets of tilted images, tilted angles, and corresponding horizontally corrected labels. Then it leveraged a mesh-to-flow transformation network to automatically correct the 2D in-plane tilt (roll) in high content fidelity without an angle prior.
@@ 3.1 Problem Formulation
Thin-Plate Spline TPS is used to simulate 2D deformation specified by a set of source control points ($\bm{P}=[p_{1},p_{2},\cdots,p_{N}]$) and the corresponding target ones ($\bm{Q}=[q_{1},q_{2},\cdots,q_{N}]$). To approximate such a non-linear transformation with minimum non-rigid distortion, it is formulated as:
@@
where $\mathcal{T}(\cdot)$ represents the desired trasformation. Then we can reach a spatial mapping function according to the derivation in [1] as follows:
@@
where $p$ is an arbitrary point on the source image and $O(r)=r^{2}logr^{2}$ is a radial basis function that indicates the impact of each control point on $p$. $C\in\mathbb{R}^{2\times 1}$, $M\in\mathbb{R}^{2\times 2}$, and $w_{i}\in\mathbb{R}^{2\times 1}$ are the transformation parameters, which can be solved with $N$ pairs of control points and dimensional constraints [32].
@@ 3.3 Dual Transformation for Correction Consistency
To this end, we propose to replace color augmentation with graphic augmentation. Particularly, we adopt content-aware rotation [14] to randomly rotate an unknown tilted image +/-3°, ensuring the augmented image is inconsistent with the original one with unnoticeable distortions.
@@ 3.4 Semi-Supervised Learning
Method Input PSNR ($\uparrow$) SSIM ($\uparrow$) FID [37] ($\downarrow$) LPIPS [38] ($\downarrow$)
1 Rotation Tilted image & angle 11.57 0.374 34.40 0.468
2 Rotation & He et al.’s rectangling [19] Tilted image & angle 17.63 0.488 15.30 0.345
3 Rotation & Nie et al.’s rectangling [11] Tilted image & angle 19.89 0.550 13.40 0.286
4 Content-aware rotation [14] Tilted image & angle 21.69 0.646 8.51 0.212
5 DRC [10] Tilted image 21.02 0.628 7.12 0.205
6 CoupledTPS (ours) Tilted image 22.29 0.679 7.90 0.197
@@ 4.2.1 Rotation Correction
Method PSNR ($\uparrow$) SSIM ($\uparrow$) FID [37] ($\downarrow$)
1 Reference 11.30 0.325 44.47
2 He et al.’s rectangling [19] 14.70 0.378 38.19
3 Nie et al.’s rectangling [11] 21.28 0.714 21.77
4 CoupledTPS (ours) 22.09 0.764 20.02
@@ 4.2.2 Rectangling
Method LineAcc [12] ($\uparrow$) ShapeAcc [12] ($\uparrow$)
1 Shih et al. [29] 66.143 97.253
2 Tan et al. [12] 66.784 97.490
3 Zhu et al. [13] 66.825 97.491
4 CoupledTPS (ours) 66.808 97.500

### v2

@@ Abstract
Thin-plate spline (TPS) is a principal warp that allows for representing elastic, nonlinear transformation with control point motions. With the increase of control points, the warp becomes increasingly flexible but usually encounters a bottleneck caused by undesired issues, e.g., content distortion. In this paper, we explore generic applications of TPS in single-image-based warping tasks, such as rotation correction, rectangling, and portrait correction. To break this bottleneck, we propose the coupled thin-plate spline model (CoupledTPS), which iteratively couples multiple TPS with limited control points into a more flexible and powerful transformation. Concretely, we first design an iterative search to predict new control points according to the current latent condition. Then, we present the warping flow as a bridge for the coupling of different TPS transformations, effectively eliminating interpolation errors caused by multiple warps. Besides, in light of the laborious annotation cost, we develop a semi-supervised learning scheme to improve warping quality by exploiting unlabeled data. It is formulated through dual transformation between the searched control points of unlabeled data and its graphic augmentation, yielding an implicit correction consistency constraint. Finally, we collect massive unlabeled data to exhibit the benefit of our semi-supervised scheme in rotation correction. Extensive experiments demonstrate the superiority and universality of CoupledTPS over the existing state-of-the-art (SoTA) solutions for rotation correction and beyond. The code and data are available at https://github.com/nie-lang/CoupledTPS.
@@ 2.1 Rotation Correction
Rigid rotation operation inclines the whole image with a certain angle. It produces rotated rectangular boundaries, which are no longer suitable for the screen of common display devices (e.g., TVs, mobile phones, etc.). To preserve the regular boundaries, He et al. proposed content-aware rotation [14] to rotate the particular contents instead of the whole image. It detected line segments by LSD [15], forced them to rotate to a certain angle, and preserved the contents and boundaries. However, it can only preserve linear structures such as pillars, buildings, etc. When there is a demand to correct a tilted image to the horizontal condition, the specific rotation angle is necessary but unavailable. To acquire this angle prior, whether it is calibrated manually or by automated algorithms [16, 17, 18], the whole process becomes cumbersome and prone to accumulate errors.
To simplify this tedious workflow and free from the constriction of angle prior, Nie et al. proposed a new and practical computer vision task, named rotation correction [10]. This task aims to automatically correct the 2D in-plane tilt (roll) with high content fidelity (preserving contents and boundaries) without the angle prior. The corrected results have the perceptually horizontal content and regular rectangular boundaries to fit common display devices. To this end, Nie et al. leveraged a mesh-to-flow transformation network and a rotation correction dataset (DRC-D) with massive triplets of tilted images, tilted angles, and corresponding horizontally corrected labels.
@@ 3.1 Problem Formulation
Thin-Plate Spline TPS is used to simulate 2D deformation specified by a set of source control points ($\bm{P}=[p_{1},p_{2},\cdots,p_{N}]^{T}$, $p_{i}\in\mathbb{R}^{2\times 1}$) and the corresponding target ones ($\bm{Q}=[q_{1},q_{2},\cdots,q_{N}]^{T}$, $q_{i}\in\mathbb{R}^{2\times 1}$ ). To approximate such a non-linear transformation with minimum non-rigid distortion, it is formulated as:
@@
where $\mathcal{T}(\cdot)$ represents the desired transformation. Then we can reach a spatial mapping function according to the derivation in [1] as follows:
@@
where $p$ is an arbitrary point on the source image and $O(r)=r^{2}logr^{2}$ is a radial basis function that indicates the impact of each control point on $p$. $C\in\mathbb{R}^{2\times 1}$, $M\in\mathbb{R}^{2\times 2}$, and $w_{i}\in\mathbb{R}^{2\times 1}$ are the transformation parameters.
To solve these parameters, we can formulate $N$ data constraints with $N$ pairs of control point correspondences ($i.e.$, $\bm{P},\bm{Q}\in\mathbb{R}^{N\times 2}$) following Eq. 2. Besides, the dimensional constraints [1],[32] are leveraged to assist the solving process as follows:
$\sum_{i=1}^{N}w_{i}=0\hskip 5.69046pt,\hskip 5.69046pt\sum_{i=1}^{N}p_{i}w_{i}^{T}=0.$ (3)
Then, we can rewrite all constraints (including data and dimensional constraints) as a form of matrix calculation and solve the parameters as:
$\begin{bmatrix}C^{T}\\ M^{T}\\ W\end{bmatrix}=\begin{bmatrix}1_{N\times 1}&\bm{P}&K\\ 0_{1\times 1}&0_{1\times 2}&1_{1\times N}\\ 0_{2\times 1}&0_{2\times 2}&\bm{P}^{T}\end{bmatrix}^{-1}\begin{bmatrix}\bm{Q}\\ 0_{2\times 2}\\ 0_{1\times 2}\end{bmatrix},$ (4)
where $W=[w_{i},...,w_{N}]^{T}$. Each element $k_{ij}$ in $K\in\mathbb{R}^{N\times N}$ is determined by $O(\parallel p_{i}-p_{j}\parallel_{2})$.
@@ 3.3 Dual Transformation for Correction Consistency
To this end, we propose to replace color augmentation with graphic augmentation. Particularly, we adopt content-aware rotation [14] to rotate an unlabeled tilted image with +/-3° randomly. This non-rigid transformation can produce an augmented image with content tilted to a different degree from the original unlabeled input, while still maintaining regular boundaries with unnoticeable distortion.
@@ 3.4 Semi-Supervised Learning
Method Input PSNR ($\uparrow$) SSIM ($\uparrow$) FID [37] ($\downarrow$) LPIPS [38] ($\downarrow$) Time ($\downarrow$)
1 Rotation Tilted image & angle 11.57 0.374 34.40 0.468 -
2 Rotation & He et al.’s rectangling [19] Tilted image & angle 17.63 0.488 15.30 0.345 -
3 Rotation & Nie et al.’s rectangling [11] Tilted image & angle 19.89 0.550 13.40 0.286 65ms
4 Content-aware rotation [14] Tilted image & angle 21.69 0.646 8.51 0.212 -
5 DRC [10] Tilted image 21.02 0.628 7.12 0.205 46ms
6 CoupledTPS (ours) Tilted image 22.29 0.679 7.90 0.197 38ms
@@ 4.2.1 Rotation Correction
Method PSNR ($\uparrow$) SSIM ($\uparrow$) FID [37] ($\downarrow$) Time ($\downarrow$)
1 Reference 11.30 0.325 44.47 -
2 He et al.’s rectangling [19] 14.70 0.378 38.19 -
3 Nie et al.’s rectangling [11] 21.28 0.714 21.77 65ms
4 CoupledTPS (ours) 22.09 0.764 20.02 38ms
@@ 4.2.2 Rectangling
Method LineAcc [12] ($\uparrow$) ShapeAcc [12] ($\uparrow$) Time ($\downarrow$)
1 Shih et al. [29] 66.143 97.253 -
2 Tan et al. [12] 66.784 97.490 -
3 Zhu et al. [13] 66.825 97.491 687ms
4 CoupledTPS (ours) 66.808 97.500 720ms
@@ 4.2.3 Wide-Angle Portrait Correction
Inference Time The comparison of inference time is also shown in Table IIIIIIV. To ensure experimental fairness, we only demonstrate the time of the methods that have released the official implementations. All the methods are tested on one RTX 3090 GPU. The inference time of our CoupledTPS may vary according to different resolutions in different tasks.
@@ 4.4 Ablation Study
TABLE VIII: Quantitative comparisons of different backbones on the DRC-D dataset.
Backbone PSNR SSIM Params Time
1 ResNet18 22.29 0.679 116MB 38ms
2 ResNet34 22.17 0.671 137MB 40ms
3 ResNet50 22.27 0.678 152MB 41ms
4 ResNet101 22.24 0.676 225MB 51ms
5 ViT-B/32 22.28 0.677 452MB 36ms
6 ViT-L/32 22.41 0.685 690MB 46ms
7 Swin-S 22.47 0.686 236MB 56ms
8 Swin-B 22.41 0.683 338MB 61ms
Backbone. We also conduct experiments to study the impact of diverse backbones on our performance. The results are illustrated in Table VIII. In the series of ResNet [34], different convolutional layers (from 18 to 101) achieve similar performance, which implies the number of convolutional layers has a limited impact on performance. Besides, we also validate our CoupledTPS on Transformer-based backbones [46] (i.e., ViT [47] and Swin Transformer [48]). It shows that our CoupledTPS can produce better results on Transformer-based backbones. Interestingly, in ViT, the larger network (ViT-L/32) can produce better performance, while it is exactly the opposite in Swin Transformer. Considering the increase in computational complexity brought by transformers, we still chose ResNet18 as our backbone in this work.